# Totally odd immersions of complete graphs in graph products

### v1

@@ Abstract
The counterexamples that Catlin used to disprove Hajós’ conjecture (that every graph $G$ contains a subdivision of $K_{\chi(G)}$) are the lexicographic product of cliques and cycles. In 1989, Lescure and Meyniel made a conjecture that is a weakening of Hajós’ and that remains open: every graph contains an immersion of $K_{\chi(G)}$. Can minimal counterexamples to this conjecture be produced through graph products? Collins, Heenehan, and McDonald recently gave a negative answer to this question for the lexicographic and Cartesian product. We consider a strengthening of the Lescure and Meyniel conjecture, different from that of Hajós, based on the notion of totally odd immersions. We study the largest totally odd immersion appearing in the graph product of two graphs. Our results imply that no minimal counterexample to this strengthened conjecture can be obtained from the Cartesian, lexicographic, direct (tensor) or strong product of graphs.
@@ 1 Introduction
Extending what at the time was the Four Color Conjecture, Hajós [12] conjectured that every graph $G$ contains a subdivision of a complete graph on $\chi(G)$ vertices (where $\chi(G)$ denotes the chromatic number of $G$). This conjecture was disproved by Catlin [3], who showed that it fails whenever $\chi(G)\geq 7$. Catlin’s counterexamples are remarkably simple, and in particular they can be seen as a particular type of graph product (lexicographic product) between a cycle and a complete graph.
When considering other graph orderings other than subdivision, one finds related (open) conjectures. If we consider the minor order, we find famous Hadwiger’s conjecture [11] which asserts that every graph $G$ contains $K_{\chi(G)}$ as a minor. Another similar conjecture is related to the immersion order, which we now define. To split off a pair of adjacent edges $uv,vw$ (with $u,v,w$ distinct) is to delete both these edges and to add the edge $uw$, if it is not already present. A graph $G$ is said to contain another graph $H$ as an immersion if $H$ can be obtained from $G$ by iteratively splitting off edges and deleting vertices and edges. The following conjecture of Lescure and Meyniel [14] has received considerable attention in recent years [1, 2, 7, 6, 8, 9, 15].
Conjecture 1.1 (Lescure and Meyniel [14]).
Every graph $G$ contains $K_{\chi(G)}$ as an immersion.
As it will be made clearer below if $G$ contains $H$ as a subdivision, then it contains $H$ as an immersion. Thus Conjecture 1.1 is a weakening of that of Hajós. Therefore it is natural to wonder whether Catlin’s counterexamples to Hajós’ conjecture can be counterexamples to Conjecture 1.1. Guyer and McDonald proved that this is not the case [10] (see also [16]), but one could still wonder whether minimal counterexamples to Conjecture 1.1 can be obtained from lexicographic products. Partially motivated by this, Collins, Heenehan and McDonald [5] studied the behaviour of complete graph immersions when one of the four standard graph products -- Cartesian, direct (or tensor), lexicographic, and strong product -- are applied. Letting $\mathrm{im}(G)$ be the order of the largest complete graph immersion of $G$, they raised the following question.
@@
Our contributions are inspired by this line of research and are concerned with a strengthening of Conjecture 1.1, different from Hajós’ conjecture, which we now introduce.
@@ 1.1 Totally odd immersions
The class of all graphs excluding a fixed graph as a immersion (or even as a subdivision) is a sparse class. On the other hand, it not hard to see that the class of complete bipartite graphs excludes $K_{3}$ as a totally odd immersion. Therefore, the following conjecture (which is inspired by one of Churchley [4]) extends Conjecture 1.1 to arbitrarily dense graph classes.
@@
Every graph $G$ contains $K_{\chi(G)}$ as a totally odd strong immersion.
While Hajós’ conjecture and Conjecture 1.3 are incomparable it is interesting to check whether Catlin’s counterexamples to Hajós’ conjecture could be counterexamples to Conjecture 1.3. This possibility was discarded by Jiménez, Quiroz and Thraves Caro [13]. Yet it is natural to wonder whether graph products can produce minimal counterexamples to Conjecture 1.3. With this as a motivation, we consider a question analogous to Question 1.2, but for the order of the largest complete graph that $G$ contains as a totally odd strong immersion, which we denote by $\mathrm{toi}(G)$. The answer to this question is the following.
@@
In Sections 2, 3 and 4 we show that Theorem 1.4 holds for the case of the strong, Cartesian and direct product, respectively. The method used in [5] already gives the case of lexicographic product for Theorem 1.4. Our results imply that no minimal counterexample Conjecture 1.3 can be obtained from the Cartesian, lexicographic, direct (tensor) or strong product of graphs. In the rest of this section we provide the definitions of the graph products which are the object of our study.
All graphs in this paper are finite, simple and loopless. Let $G$ and $H$ be graphs. The lexicographic product graph between $G$ and $H$ is the graph on vertices $V(G)\times V(H)$ and where two vertices $(g_{1},h_{1})$, $(g_{2},h_{2})$ are adjacent if $g_{1}g_{2}\in E(G)$, or $g_{1}=g_{2}$ and $h_{1}h_{2}\in E(H)$. The resulting graph is denoted $G\Circle H$. The Cartesian product between $G$ and $H$, denoted by $G\Square H$, is the graph on vertices $V(G)\times V(H)$ where two vertices $(g_{1},h_{1}),(g_{2},h_{2})$ are adjacent if $g_{1}=g_{2}$ and $h_{1}h_{2}\in E(H)$, or $h_{1}=h_{2}$ and $g_{1}g_{2}\in E(G)$. The direct product (also known as tensor product) between $G$ and $H$, denoted by $G\times H$, is the graph with vertex set $V(G)\times V(H)$ and where two vertices $(g_{1},h_{1}),(g_{2},h_{2})$ are adjacent if $g_{1}g_{2}\in E(G)$ and $h_{1}h_{2}\in E(H)$. The strong product between $G$ and $H$, denoted by $G\boxtimes H$, corresponds to the union $(G\Square H)\cup(G\times H)$. That is, $V(G\boxtimes H)=V(G)\times V(H)$ and two vertices $(g_{1},h_{1}),(g_{2},h_{2})$ are adjacent if $g_{1}=g_{2}$ and $h_{1}h_{2}\in E(H)$, or $h_{1}=h_{2}$ and $g_{1}g_{2}\in E(G)$, or $g_{1}g_{2}\in E(G)$ and $h_{1}h_{2}\in E(H)$.
2 Strong Product
In this section we study the behaviour of the parameter $\mathrm{toi}$ under the operation of strong product. We show the strong product case of Theorem 1.4, which is equivalent to proving the following theorem. Since $\chi(G\boxtimes H)\leq\chi(G)\chi(H)$ for every pair of graphs $G$ and $H$, this implies that no minimal counterexample to Conjecture 1.3 is the strong product of two graphs.
@@ 2 Strong Product
Let $G$ and $H$ be two graphs with $\mathrm{toi}(G)=t\geq 2$ and $\mathrm{toi}(H)=s\geq 2$, then $\mathrm{toi}(G\boxtimes H)\geq ts$.
@@
Let $u_{1},u_{2},\dots u_{t}$ be the terminals of the totally odd strong immersion of $K_{t}$ in $G$ and, for $1\leq i<i^{\prime}\leq t$, let $P_{ii^{\prime}}$ be the corresponding odd path joining $u_{i}$ and $u_{i^{\prime}}$. Analogously, we let $v_{1},v_{2},\dots v_{s}$ be the terminals of the totally odd strong immersion of $K_{s}$ in $G$ and $Q_{jj^{\prime}}$ be the corresponding odd path joining $v_{j}$ and $v_{j^{\prime}}$, where $1\leq j<j^{\prime}\leq s$. In the following, we show the existence of a totally odd strong immersion of $K_{ts}$ in $G\boxtimes H$ with set of terminals $\{(u_{i},v_{j})\mid 1\leq i\leq t,1\leq j\leq s\}$.
For $1\leq i<i^{\prime}\leq t$ and $1\leq j<j^{\prime}\leq s$, we let
$P_{ii^{\prime}}=u_{i}-a_{1}-a_{2}-\dots-a_{k}-u_{i^{\prime}}$
let
$Q_{jj^{\prime}}=v_{j}-b_{1}-b_{2}-\dots-b_{\ell}-v_{j^{\prime}}.$
Since $P_{ii^{\prime}}$, $Q_{jj^{\prime}}$ are odd paths, we have that $k$ and $\ell$ are even numbers.
Regarding the totally odd strong immersion of $K_{ts}$ in $G\boxtimes H$, we first construct odd paths $M_{(i,j)-(i^{\prime},j^{\prime})}$ in $G\boxtimes H$ joining the terminals $(u_{i},v_{j})$ and $(u_{i^{\prime}},v_{j^{\prime}})$, and then show that these paths are pairwise edge-disjoint.
First, whenever $(u_{i},v_{j})$ is adjacent to $(u_{i^{\prime}},v_{j^{\prime}})$ in $G\boxtimes H$, we let the path $M_{(i,j)-(i^{\prime},j^{\prime})}$ be the edge $(u_{i},v_{j})-(u_{i^{\prime}},v_{j^{\prime}})$. In any other case, we use the immersion paths in $G$ and $H$ to get the suitable paths as mentioned below. Let $1\leq i\leq t$ and $1\leq j<j^{\prime}\leq s$. We let $M_{(i,j)-(i,j^{\prime})}$ be the odd path from $(u_{i},v_{j})$ to $(u_{i},v_{j^{\prime}})$, where the first coordinates of the internal vertices are $u_{i}$ and the second coordinates moves according to the odd path $Q_{jj^{\prime}}$. More specifically saying $M_{(i,j)-(i,j^{\prime})}$ is the odd path
$(u_{i},v_{j})-(u_{i},b_{1})-(u_{i},b_{2})-\dots-(u_{i},b_{\ell})-(u_{i},b_{j^{\prime}})$
in $G\boxtimes H$. Analogously, for the case that $1\leq i<i^{\prime}\leq t$ and $1\leq j\leq s$, we let $M_{(i,j)-(i^{\prime},j)}$ be the path
$(u_{i},v_{j})-(a_{1},v_{j})-(a_{2},v_{j})-\dots(a_{k},v_{j})-(u_{i^{\prime}},v_{j}).$
Notice that all these paths are pairwise edge-disjoint.
Now, we let $1\leq i<i^{\prime}\leq t$, $1\leq j<j^{\prime}\leq s$ and focus on the description of the odd paths $M_{(i,j)-(i^{\prime},j^{\prime})}$ and $M_{(i,j^{\prime})-(i^{\prime},j)}$, the ones joining $(u_{i},v_{j})$ to $(u_{i^{\prime}},v_{j^{\prime}})$ and $(u_{i},v_{j^{\prime}})$ to $(u_{i^{\prime}},v_{j})$, respectively. Note that, when both $k$ and $\ell$ are zero, the vertices $(u_{i},v_{j})$ and $(u_{i^{\prime}},v_{j^{\prime}})$ are adjacent and hence we have defined $M_{(i,j)-(i^{\prime},j^{\prime})}$ and $M_{(i,j^{\prime})-(i^{\prime},j)}$ to be the edges $(u_{i},v_{j})-(u_{i^{\prime}},v_{j^{\prime}})$ and $(u_{i},v_{j^{\prime}})-(u_{i^{\prime}},v_{j}^{\prime})$, as we have mentioned earlier. Hence, we may assume that at least one of the numbers $k$ or $\ell$ is non-zero. Now we first define $M_{(i,j)-(i^{\prime},j^{\prime})}$ depending on $k$ and $\ell$.
If $0<k<\ell$, we define $M_{(i,j)-(i^{\prime},j^{\prime})}$ through the sequence of vertices
@@
where from $(a_{k},b_{k})$ to $(u_{i^{\prime}},v_{j^{\prime}})$ the first coordinate alternates between $a_{k}$ and $u_{i^{\prime}}$. Note that this is a valid path since $k$ and $\ell$ are of same parity.
If $0<\ell<k$, we define $M_{(i,j)-(i^{\prime},j^{\prime})}$ as
@@
where from $(a_{\ell},b_{\ell})$ to $(u_{i^{\prime}},v_{j^{\prime}})$ the second coordinate alternates between $b_{\ell}$ and $v_{j^{\prime}}$.
Otherwise, either $0\in\{k,\ell\}$ or $k=\ell$ and in those cases we define $M_{(i,j)-(i^{\prime},j^{\prime})}$ according to the rules stated in Table 1.
@@
$(u_{i},v_{j})-(u_{i^{\prime}},b_{1})-(u_{i},b_{2})-\dots-(u_{i},b_{k})-(u_{i^{\prime}},v_{j^{\prime}})$ $k=0$, $\ell>1$
$(u_{i},v_{j})-(a_{1},v_{j^{\prime}})-(a_{2},v_{j})-\dots(a_{k},v_{j})-(u_{i^{\prime}},v_{j^{\prime}})$ $\ell=0$, $k>1$
@@
Table 1: Rules for the definition of odd paths joining $(u_{i},v_{j})$ to $(u_{i^{\prime}},v_{j^{\prime}})$ in the case that $1\leq i<i^{\prime}\leq t$ and $1\leq j<j^{\prime}\leq s$.
Now we move foreword to define the odd paths $M_{(i,j^{\prime})-(i^{\prime},j)}$ joining $(u_{i},v_{j^{\prime}})$ to $(u_{i^{\prime}},v_{j})$, where $1\leq i<i^{\prime}\leq t$ and $1\leq j<j^{\prime}\leq s$. Recall that we have assumed that either $k$ or $\ell$ is a non-zero number.
If $0<k<\ell$, the following sequence of vertices defines the odd path $M_{(i,j^{\prime})-(i^{\prime},j)}$
$(u_{0},b_{j^{\prime}})-(a_{1},b_{\ell})-(a_{2},b_{\ell-1})-\dots-(a_{k},b_{\ell-k+1})-(u_{i^{\prime}},b_{\ell-k})-\dots-(a_{k},b_{1})-(u_{i^{\prime}},v_{j}),$
where from $(a_{k},b_{\ell-k+1})$ to $(u_{i^{\prime}},v_{j})$ the first coordinate alternates between $a_{k}$ and $u_{i^{\prime}}$.
If $0<\ell<k$, then we define $M_{(i,j^{\prime})-(i^{\prime},j)}$ as
$(u_{i},b_{j^{\prime}})-(a_{1},b_{\ell})-(a_{2},b_{\ell-1})-\dots-(a_{\ell},b_{1})-(a_{\ell+1},v_{j})-(a_{\ell+2},b_{1})-\dots-(a_{k},b_{1})-(u_{i^{\prime}},v_{j}),$
where from $(a_{\ell},b_{1})$ to $(a_{k+1},b_{0})$ the second coordinate alternates between $b_{1}$ and $u_{i}$.
In the rest of the cases (either $k=\ell$ or $0\in\{k,\ell\}$), we define $M_{(i,j^{\prime})-(i^{\prime},j)}$ according to the rules stated in Table 2.
definition of $M_{(i,j)-(i^{\prime},j^{\prime})}$ condition
$(u_{i},v_{j^{\prime}})-(a_{1},b_{\ell})-(a_{2},b_{\ell-1})-\dots-(a_{k},b_{1})-(u_{i^{\prime}},v_{j})$. $k=\ell$
$(u_{i},v_{j^{\prime}})-(u_{i^{\prime}},b_{\ell})-(u_{i},b_{\ell-1})-(u_{i^{\prime}},b_{\ell-2})-\dots-(u_{i^{\prime}},v_{j})$ $k=0$, $\ell>1$
$(u_{i},v_{j^{\prime}})-(a_{1},v_{j})-(a_{2},v_{j^{\prime}})-(a_{3},v_{j})-\dots-(u_{i^{\prime}},v_{j})$ $\ell=0$, $k>1$

### v4

@@ Abstract
For a graph $G$, let $\mathrm{im}(G)$ denote the maximum integer $t$ such that $G$ contains $K_{t}$ as an immersion. A recent paper of Collins, Heenehan, and McDonald (2023) studied the behaviour of this parameter under graph products, asking how large can $\mathrm{im}(G\ast H)$ be in terms of $\mathrm{im}(G)$ and $\mathrm{im}(H)$, when $\ast$ is one of the four standard graph products. We consider a similar question for the parameter $\mathrm{toi}(G)$ which denotes the maximum integer $t$ such that $G$ contains $K_{t}$ as a totally odd immersion. As an application, we obtain that no minimum counterexample to the immersion-analogue of the Odd Hadwiger Conjecture can be obtained from the Cartesian, direct (tensor), lexicographic or strong product of graphs.
@@ 1 Introduction
To split off a pair of adjacent edges $uv,vw$ (with $u,v,w$ distinct) is to delete both these edges and to add the edge $uw$, if it is not already present. A graph $G$ is said to contain another graph $H$ as an immersion if $H$ can be obtained from $G$ by iteratively splitting off edges and deleting vertices and edges. In 1965, Nash-Williams conjectured [18] that graphs are well-quasi-ordered by the immersion relation, and his conjecture was confirmed [19] by Robertson and Seymour in 2010. By $\mathrm{im}(G)$ we denote the maximum integer $t$ such that $G$ contains an immersion of $K_{t}$. This is a well-studied parameter. DeVos, McDonald, Mohar and Scheide [8], and, independently, Wollan [20] gave a structure theorem for graphs $G$ with $\mathrm{im}(G)<t$. Liu [17] provided a different structure theorem for excluded immersions. Moreover, the following conjecture, which relates $\mathrm{im}(G)$ to the chromatic number, $\chi(G)$, of $G$ has received considerable attention in recent years [1, 2, 7, 6, 9, 10, 16].
Conjecture 1.1 (Abu-Khzam and Langston [15]).
For every graph $G$ we have $\chi(G)\leq\mathrm{im}(G)$.
In a recent paper, Collins, Heenehan and McDonald [5] studied the behaviour of complete graph immersions when one of the four standard graph products - Cartesian, direct (or tensor), lexicographic, and strong product - are applied. They raised the following question.
@@
Our contributions are inspired by this line of research and are concerned with a restriction of the immersion relation which we now introduce.
@@ 1.1 Totally odd immersions
Churchley [4] gave a structure theorem for graphs excluding $K_{t}$ as a totally odd immersion. Let $\mathrm{toi}(G)$ denote the maximum $t$ such that $G$ contains a totally odd strong immersion of $K_{t}$. While the class of all graphs $G$ with $\mathrm{im}(G)\leq t$ is a sparse class (one with bounded degeneracy), it is not hard to see that if $B$ is a complete bipartite graph, then $\mathrm{toi}(B)\leq 2$. Therefore, the following conjecture (which is inspired by one of Churchley [4]) extends Conjecture 1.1 to arbitrarily dense graph classes.
@@
For every graph $G$, we have $\chi(G)\leq\mathrm{toi}(G)$.
Extending the study of totally odd immersions, we consider a question analogous to Question 1.2 for the parameter $\mathrm{toi}$. The answer to this question is the following.
@@
In Sections 2 and 3 we show that Theorem 1.4 holds for the case of the direct and the cartesian product, respectively. The proof of Theorem 3 in [5] already proves Theorem 1.4 for the lexicographic product. Similarly, of Theorem 6 in [21] gives Theorem 1.4 for the strong product.
We apply Theorem 1.4 to give evidence of Conjecture 1.3 as follows. A strengthening of Conjecture 1.1, different from Conjecture 1.3, had been proposed earlier by Hajós [13], who conjectured that every graph $G$ contains a subdivision of a complete graph on $\chi(G)$ vertices. Hajós’ conjecture was disproved by Catlin [3], who showed that it fails whenever $\chi(G)\geq 7$. Catlin’s counterexamples are remarkably simple, and in particular they can be seen as a particular type of graph product (lexicographic product) between a cycle and a complete graph. While Hajós’ conjecture and Conjecture 1.3 are incomparable it is interesting to check whether Catlin’s counterexamples to Hajós’ conjecture could be counterexamples to Conjecture 1.3. This possibility was discarded by Jiménez, Quiroz and Thraves Caro [14]. Yet it is natural to wonder whether graph products can produce minimal counterexamples to Conjecture 1.3. This is, however, discarded by the following.
Corollary 1.5.
Let $\ast$ be any of the four standard graph products. No minimum counterexample to Conjecture 1.3 is of the form $G\ast H$.
We prove now that this is the case for the lexicographic product; the same proof works for the strong product. For the other products, the reader can check the corresponding sections of this paper. As mentioned earlier, by the proof of [5, Theorem 3] we have that $\mathrm{toi}(G\circ H)\geq\mathrm{toi}(G)\mathrm{toi}(H)$ for every $G$ and $H$. So suppose for a contradiction that there are graphs $G$ and $H$ so that $G\circ H$ is a minimal counterexample to Conjecture 1.3. By minimality we have $\chi(G)\leq\mathrm{toi}(G)$ and $\chi(H)\leq\mathrm{toi}(H)$. It is well known that $\chi(G\circ H)\leq\chi(G)\chi(H)$, and so we obtain, $\chi(G\circ H)\leq\mathrm{toi}(G\circ H)$, a contradiction.
We end this section with the definitions of the graph products which are the object of our study.
All graphs in this paper are finite, simple and loopless. Let $G$ and $H$ be graphs. The lexicographic product graph between $G$ and $H$ is the graph on vertices $V(G)\circ V(H)$ and where two vertices $(g_{1},h_{1})$, $(g_{2},h_{2})$ are adjacent if $g_{1}g_{2}\in E(G)$, or $g_{1}=g_{2}$ and $h_{1}h_{2}\in E(H)$. The resulting graph is denoted $G\circ H$. The Cartesian product between $G$ and $H$, denoted by $G\Square H$, is the graph on vertices $V(G)\times V(H)$ where two vertices $(g_{1},h_{1}),(g_{2},h_{2})$ are adjacent if $g_{1}=g_{2}$ and $h_{1}h_{2}\in E(H)$, or $h_{1}=h_{2}$ and $g_{1}g_{2}\in E(G)$. The direct product (also known as tensor product) between $G$ and $H$, denoted by $G\times H$, is the graph with vertex set $V(G)\times V(H)$ and where two vertices $(g_{1},h_{1}),(g_{2},h_{2})$ are adjacent if $g_{1}g_{2}\in E(G)$ and $h_{1}h_{2}\in E(H)$. The strong product between $G$ and $H$, denoted by $G\boxtimes H$, corresponds to the union $(G\Square H)\cup(G\times H)$. That is, $V(G\boxtimes H)=V(G)\times V(H)$ and two vertices $(g_{1},h_{1}),(g_{2},h_{2})$ are adjacent if $g_{1}=g_{2}$ and $h_{1}h_{2}\in E(H)$, or $h_{1}=h_{2}$ and $g_{1}g_{2}\in E(G)$, or $g_{1}g_{2}\in E(G)$ and $h_{1}h_{2}\in E(H)$.
2 Direct Product
In this section we study the behaviour of the parameter $\mathrm{toi}$ under the operation of direct product. The direct product of a bipartite graph with any other graph is a bipartite graph. Thus, $\mathrm{toi}(G\times H)\leq 2$, if either $G$ or $H$ is bipartite. Thus, we only need to study the case of $\mathrm{toi}(G\times H)$ when neither $G$ nor $H$ is a bipartite graph.
@@ 2 Direct Product
$\mathrm{toi}(G\times H)\geq\mathrm{toi}(K_{t}\times K_{s}).$
@@
Let $u_{1},u_{2},\dots u_{t}$ be the terminals of a totally odd strong immersion of $K_{t}$ in $G$ and, for $1\leq i<i^{\prime}\leq t$, let $P_{ii^{\prime}}$ be the corresponding odd path joining $u_{i}$ and $u_{i^{\prime}}$. Analogously, we let $v_{1},v_{2},\dots v_{s}$ be the terminals of a totally odd strong immersion of $K_{s}$ in $H$ and $Q_{jj^{\prime}}$ be the corresponding odd path joining $v_{j}$ and $v_{j^{\prime}}$, where $1\leq j<j^{\prime}\leq s$. Suppose that $\mathrm{toi}(K_{t}\times K_{s})=r$, and hence, $K_{t}\times K_{s}$ has a totally odd immersion of $K_{r}$. Let $V(K_{t})=\{x_{1},\ldots,x_{t}\}$, $V(K_{s})=\{y_{1},\ldots,y_{s}\}$, and $c_{1},\cdots,c_{r}\subseteq\{(x_{i},y_{j})\mid 1\leq i\leq t,1\leq j\leq s\}$ be the terminals of a totally odd immersion of $K_{r}$ in $K_{t}\times K_{s}$. Let us consider the injective function $\phi:K_{t}\times K_{s}\rightarrow G\times H$ as $\phi(x_{i},y_{j})=(u_{i},v_{j})$, for all $1\leq i\leq t$ and $1\leq j\leq s$. Our plan is to construct a totally odd strong immersion of $K_{r}$ in $G\times H$ with set of terminals
$\{\phi(c)\in\{(u_{i},v_{j})\mid 1\leq i\leq t,1\leq j\leq s\}\mid c\in\{c_{1},\ldots,c_{r}\}\}.$
For $1\leq i<i^{\prime}\leq t$ and $1\leq j<j^{\prime}\leq s$, we let $P_{ii^{\prime}}=u_{i}-a_{1}-a_{2}-\dots-a_{k}-u_{i^{\prime}}$ and $Q_{jj^{\prime}}=v_{j}-b_{1}-b_{2}-\dots-b_{\ell}-v_{j^{\prime}}$, as before. Since $P_{ii^{\prime}}$, $Q_{jj^{\prime}}$ are odd paths, $k$ and $\ell$ are even numbers. We are going to define two odd paths $M_{(i,j)-(i^{\prime},j^{\prime})}$ and $M_{(i,j^{\prime})-(i^{\prime},j)}$ in $G\times H$, connecting $(u_{i},v_{j})$ to $(u_{i^{\prime}},v_{j^{\prime}})$ and $(u_{i},v_{j^{\prime}})$ to $(u_{i^{\prime}},v_{j})$, respectively, as follows.
When both $k$ and $\ell$ are zero, the vertices $(u_{i},v_{j})$ and $(u_{i^{\prime}},v_{j^{\prime}})$ are adjacent, and hence $M_{(i,j)-(i^{\prime},j^{\prime})}$ and $M_{(i,j^{\prime})-(i^{\prime},j)}$ are defined to be the edges $(u_{i},v_{j})-(u_{i^{\prime}},v_{j^{\prime}})$ and $(u_{i},v_{j^{\prime}})-(u_{i^{\prime}},v_{j}^{\prime})$. Hence, we may assume that at least one of the numbers $k$ or $\ell$ is at least one. We split the definition of the paths into cases depending on the relation between $k$ and $\ell$.
Case: $0<k<\ell$. We define $M_{(i,j)-(i^{\prime},j^{\prime})}$ through the sequence of vertices
@@
where from $(a_{k},b_{k})$ to $(u_{i^{\prime}},v_{j^{\prime}})$ the first coordinate alternates between $a_{k}$ and $u_{i^{\prime}}$, and the second coordinate continues moving along the path $Q_{jj^{\prime}}$. Note that this is a valid path since $k$ and $\ell$ are of same parity. Now, we define $M_{(i,j^{\prime})-(i^{\prime},j)}$ through the sequence of vertices
$(u_{i},v_{j^{\prime}})-(a_{1},b_{\ell})-(a_{2},b_{\ell-1})-\dots-(a_{k},b_{\ell-k+1})-(u_{i^{\prime}},b_{\ell-k})-\dots-(a_{k},b_{1})-(u_{i^{\prime}},v_{j}),$
where from $(a_{k},b_{\ell-k+1})$ to $(u_{i^{\prime}},v_{j})$ the first coordinate alternates between $a_{k}$ and $u_{i^{\prime}}$, and the second coordinate continues moving along the path $Q_{jj^{\prime}}$ from $v_{j^{\prime}}$ to $v_{j}$.
We argue that paths $M_{(i,j)-(i^{\prime},j^{\prime})}$ and $M_{(i,j^{\prime})-(i^{\prime},j)}$ are edge-disjoint. Actually, the paths are vertex-disjoint. In $M_{(i,j)-(i^{\prime},j^{\prime})}$ there are five types of vertices: the end vertices $(u_{i},v_{j})$ and $(u_{i^{\prime}},v_{j^{\prime}})$, vertices $(a_{p},b_{p})$, vertices $(a_{k},b_{even})$, and vertices $(u_{i^{\prime}},b_{odd})$. None of these vertices appear in $M_{(i,j^{\prime})-(i^{\prime},j)}$, since vertices in $M_{(i,j^{\prime})-(i^{\prime},j)}$ are of the following six types: the end vertices $(u_{i},v_{j^{\prime}})$ and $(u_{i^{\prime}},v_{j})$, vertices $(a_{odd},b_{even})$, $(a_{even},b_{odd})$, vertices $(a_{k},b_{odd})$, and vertices $(u_{i^{\prime}},b_{even})$.
Case: $0<\ell<k$. We define $M_{(i,j)-(i^{\prime},j^{\prime})}$ through the sequence of vertices
@@
where from $(a_{\ell},b_{\ell})$ to $(u_{i^{\prime}},v_{j^{\prime}})$ the second coordinate alternates between $b_{\ell}$ and $v_{j^{\prime}}$, and the first coordinate continues moving along the path $P_{ii^{\prime}}$. Now, we define $M_{(i,j^{\prime})-(i^{\prime},j)}$ through the sequence of vertices
$(u_{i},v_{j^{\prime}})-(a_{1},b_{\ell})-(a_{2},b_{\ell-1})-\dots-(a_{\ell},b_{1})-(a_{\ell+1},v_{j})-(a_{\ell+2},b_{1})-\dots-(a_{k},b_{1})-(u_{i^{\prime}},v_{j}),$
where from $(a_{\ell},b_{1})$ to $(u_{i^{\prime}},v_{j})$ the second coordinate alternates between $b_{1}$ and $v_{j}$, and the first coordinate continues moving along the path $P_{ii^{\prime}}$.
As in the previous case, we can argue that paths $M_{(i,j)-(i^{\prime},j^{\prime})}$ and $M_{(i,j^{\prime})-(i^{\prime},j)}$ are vertex-disjoint. In $M_{(i,j)-(i^{\prime},j^{\prime})}$ there are five types of vertices: the end vertices $(u_{i},v_{j})$ and $(u_{i^{\prime}},v_{j^{\prime}})$, vertices $(a_{p},b_{p})$, vertices $(a_{even},b_{\ell})$, and vertices $(a_{odd},v_{j^{\prime}})$. None of these vertices appear in $M_{(i,j^{\prime})-(i^{\prime},j)}$, since vertices in $M_{(i,j^{\prime})-(i^{\prime},j)}$ are of the following six types: the end vertices $(u_{i},v_{j^{\prime}})$ and $(u_{i^{\prime}},v_{j})$, vertices $(a_{odd},b_{even})$, $(a_{even},b_{odd})$, vertices $(a_{even},b_{1})$, and vertices $(a_{odd},v_{j})$.
Cases: either $0\in\{k,\ell\}$ or $k=\ell$. Definitions are given according to the rules stated in Table 1 and Table 2.
@@
$(u_{i},v_{j})-(u_{i^{\prime}},b_{1})-(u_{i},b_{2})-\dots-(u_{i},b_{\ell})-(u_{i^{\prime}},v_{j^{\prime}})$ $k=0$, $\ell\geq 2$
$(u_{i},v_{j})-(a_{1},v_{j^{\prime}})-(a_{2},v_{j})-\dots-(a_{k},v_{j})-(u_{i^{\prime}},v_{j^{\prime}})$ $\ell=0$, $k\geq 2$
@@
Table 1: Rules for the definition of odd paths joining $(u_{i},v_{j})$ to $(u_{i^{\prime}},v_{j^{\prime}})$ in the case that $1\leq i<i^{\prime}\leq t$ and $1\leq j<j^{\prime}\leq s$. In the subcase $k=0,\ell\geq 2$, the first coordinate alternates between $u_{i}$ and $u_{i^{\prime}}$, and the second coordinate moves along the path $Q_{jj^{\prime}}$. In the subcase $\ell=0,k\geq 2$, the second coordinate alternates between $v_{j}$ and $v_{j^{\prime}}$, and the first coordinate moves along the path $P_{ii^{\prime}}$.
definition of $M_{(i,j^{\prime})-(i^{\prime},j)}$ condition
$(u_{i},v_{j^{\prime}})-(a_{1},b_{\ell})-(a_{2},b_{\ell-1})-\dots-(a_{k},b_{1})-(u_{i^{\prime}},v_{j})$ $k=\ell$
$(u_{i},v_{j^{\prime}})-(u_{i^{\prime}},b_{\ell})-(u_{i},b_{\ell-1})-\dots-(u_{i},v_{j})-(u_{i^{\prime}},v_{j})$ $k=0$, $\ell\geq 2$
$(u_{i},v_{j^{\prime}})-(a_{1},v_{j})-(a_{2},v_{j^{\prime}})-\dots-(a_{k},v_{j^{\prime}})-(u_{i^{\prime}},v_{j})$ $\ell=0$, $k\geq 2$